OpenStudy (anonymous):

I. Given the following polynomial: II. Use the quadratic formula to solve the following:

OpenStudy (anonymous):

help me i dont kno mathat all

Nnesha (nnesha):

what's the equation ?

Nnesha (nnesha):

post ur question here :=)

Nnesha (nnesha):

`Given the following polynomial:` ???

OpenStudy (anonymous):

look at the attachment

Nnesha (nnesha):

i didn't find the equation can you please post it here ?

OpenStudy (anonymous):

how do i do that?

Nnesha (nnesha):

like type in the equation toool ?

Nnesha (nnesha):

take a screenshot ?

OpenStudy (anonymous):

 The value of the discriminant is 169.  There are 2 real roots.  There are 2 irrational roots.  The graph intersects the y-axis twice.  The parabola is directed upward.  The axis of symmetry is located at  The vertex is located at:  The roots are:  The graph intersects the y axis at (0, -15).  The graph intersects the x-axis at (-5, 0) and (1.5, 0) II. Use the quadratic formula to solve the following:

Nnesha (nnesha):

what's the polynomial ?

Nnesha (nnesha):

there should be an equation Ax^2+Bx+C like this

OpenStudy (anonymous):

2xexponet2+7x-15=0

Nnesha (nnesha):

\[\huge\rm 2x^2+7x-15=0\] quadratic equation ax^2+bx+c quadratic equation where a=leading coefficient b=middle term c-constant term so what is a b and c in that equation ? \[\huge\rm x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]quadratic formula

Nnesha (nnesha):

\(\huge\color{reD}{\rm b^2-4ac}\) `Discriminant` you can use this to find if the equation is factorable or not if ` b^2-4ac > 0` then there are 2 real zeros if ` b^2-4ac = 0` then there is one real root if ` b^2-4ac < 0` then you will get two complex roots (no -x-intercept)

Nnesha (nnesha):

substitute a b and c value into b^2 -4ac do you get 169 ?

OpenStudy (anonymous):

idk how to domath miss

Nnesha (nnesha):

Ax^2+Bx+C standard form equation of parabola where a = leading coefficient b= middle term c=constant term now look at this equation 2x^2+7x-15 what is a b and c ?

OpenStudy (anonymous):

2x is a 7x is b

Nnesha (nnesha):

just the coefficient 2 =a 7=b c = ?

OpenStudy (anonymous):

15

Nnesha (nnesha):

-15

Nnesha (nnesha):

alright \[\huge\rm b^2-4ac\] substitute b a and c values then sovle

Nnesha (nnesha):

solve*

OpenStudy (anonymous):

how

Nnesha (nnesha):

just replace b with 7 a with 2 c with -15

Nnesha (nnesha):

here is how i would replace b with 7...... \[\rm (7)^2-4ac\] ur turn replace a with 2 and c with -15

OpenStudy (anonymous):

49-8-15 that what i got

Nnesha (nnesha):

it's 4 times a times c it would bbe \[49 -8(-15)\] now solve

Nnesha (nnesha):

hello ?

OpenStudy (anonymous):

idk

Nnesha (nnesha):

that's not an answer you don't know or you don't want to do it ? you can use calculator just put 49-8(-15) into the calculator that's it

OpenStudy (anonymous):

56

Nnesha (nnesha):

that's not correct!

Nnesha (nnesha):

49 -8(-15) first multiply -8 times -15 then add 49

OpenStudy (anonymous):

thats what i got both times

Nnesha (nnesha):

-8 times -15 = ?

OpenStudy (anonymous):

1065

Nnesha (nnesha):

did you really get that ?

OpenStudy (anonymous):

yea

Nnesha (nnesha):

guess what ... try again make sure you typed -8 times -15 into the calculator

OpenStudy (anonymous):

120

Nnesha (nnesha):

yes right 49+120 = ? that would be discriminant

OpenStudy (anonymous):

169

Nnesha (nnesha):

yes right so first sentence is right ? ` The value of the discriminant is 169`